# A Study On The Graph Formulation Of Union Closed Conjecture

### v2

@@ Abstract
The Union Closed Conjecture is one of the most renowned problems in Combinatorics. Its appeal lies in the simplicity of its statement contrasted with the potential complexity of its resolution. The conjecture posits that, in any union-closed family of sets, there exists at least one element that appears in at least half of the sets within the family.
@@
Keywords: Union Closed Conjecture, Frequency, Family of sets, Member sets, Pendant vertices.
@@ 1 Introduction
[…]
@@
    When the family includes either a singleton or a set of maximal cardinality. [6, 7]
@@
In 2013, Henning Bruhn introduced a graph-theoretic approach to the Union Closed Conjecture (UCC) [2]. A subset of vertices in a graph is termed stable if no two vertices in it are adjacent. A stable set is maximal if no additional vertices can be added without losing this property. The graph formulation asserts that for any bipartite graph with at least one edge, there exists a vertex in each bipartite class that lies in at most half of the maximal stable sets. This vertex is called rare.
In Section 2, we establish a comprehensive connection between the set-based and graph-based formulations of the conjecture. Using this connection, we present the graph version of the known results as a natural consequence. In Section 3, we introduce a theorem that confirms the validity of the UCC for graphs with pendant vertices under specific configurations, and we explore the implications of this result.
@@ 2 Connecting The Two Formulations
The union-closed family generated by a family $\mathcal{F}$, denoted $\langle\mathcal{F}\rangle$, consists of the collection of all unions of its member sets, including the empty set $\emptyset$. For a graph, the set of neighbors of a vertex subset $P$ is denoted by $N(P)$ [^1] When multiple graphs are involved, we use $N_{\scriptscriptstyle G}(P)$ to specify the graph.. If $P=\{x\}$ is a singleton, we use $N(x)$ for convenience.
We define the incidence graph of $\mathcal{F}$ as the bipartite graph with vertex set $\mathcal{F}\cup U(\mathcal{F})$ and edge set $\{(S,x):S\in\mathcal{F},x\in U(\mathcal{F}),x\in S\}$. For a bipartite graph $G(X,Y)$, we refer to the incidence family of $X$ as $\mathcal{F}^{\scriptscriptstyle X}=\{N(y):y\in Y\}$. If there are no isolated vertices [^2] Isolated vertices do not affect the conjecture, and thus we exclude them from consideration. in $X$, then its universe is given by $U(\mathcal{F}^{\scriptscriptstyle X})=X$. Throughout this section, we will refer to the two vertex classes of the bipartite graph $G$ as $X$ and $Y$.
@@
Every maximal stable set $S$ satisfies $S\cap X=X\setminus N(Y^{\prime})$ for some $Y^{\prime}\subseteq Y$. It is important to note that no two maximal stable sets $S_{\scriptscriptstyle 1}$ and $S_{\scriptscriptstyle 2}$ can have $S_{\scriptscriptstyle 1}\cap X=S_{\scriptscriptstyle 2}\cap X$. Therefore, if a vertex $x\in X$ is rare in $G$, then it is also rare in the family $\{X\setminus(N(Y^{\prime}))\hskip 5.69046pt|\hskip 5.69046ptY^{\prime}\subseteq Y\}$. Consequently, $x$ is abundant in the family $\{N(Y^{\prime})\hskip 5.69046pt|\hskip 5.69046ptY^{\prime}\subseteq Y\}$. Now, let $F=\{N(y^{\prime})\hskip 5.69046pt|\hskip 5.69046pty\in Y\}$, and observe that the universe $U(F)$ is nothing but $N(Y^{\prime})$. It is also evident that $\{N(Y^{\prime})\hskip 5.69046pt|\hskip 5.69046ptY^{\prime}\subseteq Y\}=\{U(F)\hskip 5.69046pt|\hskip 5.69046ptF\subseteq\mathcal{F}^{\scriptscriptstyle X}\}=\langle\mathcal{F}^{\scriptscriptstyle X}\rangle$. The converse direction follows similarly from the above steps. ∎
@@
  • A bipartite graph with both classes of cardinality at most 12 satisfies Frankl’s conjecture.
@@
  • If $|Y|\geq\frac{2}{3}2^{\scriptscriptstyle|X|}$, then the graph satisfies Frankl’s conjecture. [cf. Result 4]
  • If a bipartite graph $G$ contains a pendant vertex, then its neighbor is rare in $G$. Additionally, if $G$ contains a vertex of degree 2, then there exists a rare vertex in the other bipartite class. [cf. Result 5]
  • Let $\mathcal{F}$ be a family in which every member set has a maximum size of 3, and each element within the sets has a frequency that does not exceed 3. Then $\langle\mathcal{F}\rangle$ satisfies the union closed conjecture. [cf. subcubic graphs satisfies UCC [2]]
@@
  • Let $G$ be a bipartite graph with a minimum degree of $\delta$. If both bipartite classes have cardinalities at most $2\delta$, then $G$ satisfies Frankl’s conjecture.
  • An $r$-regular bipartite graph with both of its bipartite classes having cardinality at most $2r$ satisfies Frankl’s conjecture.
The following three conjectures are equivalent to the UCC, as can be demonstrated by Proposition 2.1.
Conjecture 2.2.
Any bipartite graph with no pendant vertex satisfy Frankl’s conjecture.
Conjecture 2.3.
Any bipartite graph with exactly one pendant vertex satisfy Frankl’s conjecture.
Conjecture 2.4.
Any bipartite graph with at least one pendant vertex satisfy Frankl’s conjecture.
This helps to narrow down the specific classes of graphs that need to be considered in order to demonstrate the conjecture. In the subsequent section, our focus will be on graphs that possess pendant vertices positioned in specific locations.
@@ 3 Some New Graph Results
We will see some interesting results related to the graph formulation of the conjecture. Before that let us go through some definitions and notations we might encounter here.
For a graph $G$, we denote the set of all maximal stable sets by $\mathcal{B}_{\scriptscriptstyle G}$. The cardinality of $\mathcal{B}_{\scriptscriptstyle G}$ is denoted by $w_{\scriptscriptstyle G}$, which represents the number of maximal stable sets of $G$. For any vertex $x\in V(G)$, the number of maximal stable sets containing $x$ is represented by $w_{\scriptscriptstyle G}(x)$. We use the notation $N_{\scriptscriptstyle G}(P)$ to represent the set of neighbors of a vertex subset $P$ in graph $G$. When $P$ is a singleton, say $P=\{x\}$, we will simply use $N_{\scriptscriptstyle G}(x)$.
@@
If a disconnected graph $G$ has one of its connected components satisfying Frankl’s conjecture, then $G$ will also satisfy the conjecture.
@@
As Frankl’s conjecture is about finite graphs, we only need to consider finitely many connected components. We will show this for the case of two components, and the rest of the proof follows from basic induction arguments. Suppose $G$ has two components $P$ and $Q$. Let $S_{\scriptscriptstyle 1}$ and $S_{\scriptscriptstyle 2}$ be two maximal stable sets of $P$ and $Q$ respectively. One can notice that $S_{\scriptscriptstyle 1}\cup S_{\scriptscriptstyle 2}$ is a maximal stable set of $G$. For any maximal stable set $S$ of $G$, observe that $S\cap V(P)$ and $S\cap V(Q)$ are maximal stable sets of $P$ and $Q$ respectively. This gives the number of maximal stable sets of $G$, $w_{\scriptscriptstyle G}=w_{\scriptscriptstyle P}w_{\scriptscriptstyle Q}$. Using similar arguments, one can obtain that for every $x\in V(P)$, $w_{\scriptscriptstyle G}(x)=w_{\scriptscriptstyle P}(x)w_{\scriptscriptstyle Q}$.
Let $r$ be a rare vertex in $P$. Then by definition, $w_{\scriptscriptstyle P}-2w_{\scriptscriptstyle P}(r)\geq 0$. As a result, $w_{\scriptscriptstyle G}-2w_{\scriptscriptstyle G}(r)=w_{\scriptscriptstyle P}w_{\scriptscriptstyle Q}-2w_{\scriptscriptstyle P}(r)w_{\scriptscriptstyle Q}\geq 0$ and $r$ is rare in $G$. Thus, whenever $P$ satisfies the conjecture, it guarantees that it holds for $G$ as well. ∎
We will now present a more intriguing result. A decomposition of a graph $G$ is a set of edge-disjoint subgraphs $H_{1},H_{2},\ldots,H_{n}$ such that $\bigcup_{\scriptscriptstyle i=1}^{\scriptscriptstyle n}H_{i}=G$. A vertex $v\in V(G)$ is defined as 2-layered if every vertex in $N_{\scriptscriptstyle G}(v)$ has a pendant vertex adjacent to it.
@@
Let $G$ be a bipartite graph and $\{C,H\}$ a decomposition of $G$. Suppose the vertices in $V(C)\cap V(H)$ belong to the same bipartite class of $G$ and are all 2-layered. Then, all rare vertices in $C$ or $H$ will remain rare in $G$.
Before presenting the proof of the theorem, we need to establish some notations and lemmas. The following notations will be used throughout the proof. In light of Proposition 3.1, it suffices to verify the theorem for connected graphs. Consequently, we assume without loss of generality that $G$ is connected.
Let $m\in V(C)\cap V(H)$, and suppose $N_{\scriptscriptstyle C}(m)=\phi$. In this case, we can simply remove the vertex from $C$ to get a new decomposition. Similarly, if $N_{\scriptscriptstyle H}(m)=\emptyset$, we can remove $m$ from $H$. This allows us to redefine $C$ and $H$ in such a way that every vertex in $V(C)\cap V(H)$ has neighbors in both $C$ and $H$. Let $1,2,\ldots,n$ be the vertices common to these subgraphs $C$ and $H$. We define $[n]=\{i\in\mathbb{N}\mid i\leq n\}$, and accordingly, we represent this common vertex set by $[n]$. For any subset $\Theta\subseteq[n]$, we denote its complement by $\Theta^{\mathsf{c}}=[n]\setminus\Theta$.
Recall the definitions of $w_{\scriptscriptstyle G}$ and $\mathcal{B}_{\scriptscriptstyle G}$. Here, we will slightly generalize these definitions.For any vertex subsets $P,Q\subseteq V(G)$ and any subgraph $H$ of $G$, the notation $\mathcal{B}_{\scriptscriptstyle G}(P,\overline{Q})$ denotes the set of all maximal stable sets of $G$ that include all vertices in $P$ and exclude all vertices in $Q$. Similarly, $\mathcal{B}_{\scriptscriptstyle H}(P,\overline{Q})$ represents the set of all maximal stable sets of $H$ with the same inclusion and exclusion criteria. The cardinality of $\mathcal{B}_{\scriptscriptstyle G}(P,\overline{Q})$ and $\mathcal{B}_{\scriptscriptstyle H}(P,\overline{Q})$ is denoted by $w_{\scriptscriptstyle G}(P,\overline{Q})$ and $w_{\scriptscriptstyle H}(P,\overline{Q})$. This notation is somewhat informal but helps to reduce the complexity of larger notations.
@@
For any $\Theta\subseteq[n]$, $\Gamma\subseteq\Theta^{\mathsf{c}}$, and $b\in V(C\setminus\Gamma)$, the following identities hold:
$w_{\scriptscriptstyle C\setminus\Gamma}(\Theta,\overline{\Theta^{\mathsf{c}}\cup N_{\scriptscriptstyle C}(\Gamma)})=w_{\scriptscriptstyle C}(\Theta\cup\Gamma,\overline{\Theta^{\mathsf{c}}\setminus\Gamma})$ (1)
$w_{\scriptscriptstyle C\setminus\Gamma}(\{b\}\cup\Theta,\overline{\Theta^{\mathsf{c}}\cup N_{\scriptscriptstyle C}(\Gamma)})=w_{\scriptscriptstyle C}(\{b\}\cup\Theta\cup\Gamma,\overline{\Theta^{\mathsf{c}}\setminus\Gamma})$ (2)
@@
Consider the mapping $f:\mathcal{B}_{\scriptscriptstyle C\setminus\Gamma}(\Theta,\overline{\Theta^{\mathsf{c}}\cup N_{\scriptscriptstyle C}(\Gamma)})\to\mathcal{B}_{\scriptscriptstyle C}(\Theta\cup\Gamma,\overline{\Theta^{\mathsf{c}}\setminus\Gamma})$ defined by $f(B)=B\cup\Gamma$. Let $B\in\mathcal{B}{\scriptscriptstyle C\setminus\Gamma}(\Theta,\overline{\Theta^{\mathsf{c}}\cup N{\scriptscriptstyle C}(\Gamma)})$, observe that $f(B)$ includes all vertices from $\Theta\cup\Gamma$ and excludes all vertices from $\Theta^{\mathsf{c}}\setminus\Gamma$. Since $B$ is a stable set in $C\setminus\Gamma$ and contains no vertices from $N_{\scriptscriptstyle C}(\Gamma)$, it follows that $f(B)$ is stable in $C$. Assume, for contradiction, that $f(B)$ is not a maximal stable set in $C$. Then, there exists a vertex $v\in V(C)$ such that $v\notin f(B)$ and $B\cup\Gamma\cup\{v\}$ is stable in $C$. This implies that $B\cup\{v\}$ is stable in $C\setminus\Gamma$, contradicting the maximality of $B$ in $C\setminus\Gamma$. Hence, $f(B)\in\mathcal{B}_{\scriptscriptstyle C}(\Theta\cup\Gamma,\overline{\Theta^{\mathsf{c}}\setminus\Gamma})$.
@@
Next, consider a set $D\in\mathcal{B}_{\scriptscriptstyle C}(\Theta\cup\Gamma,\overline{\Theta^{\mathsf{c}}\setminus\Gamma})$. We define the inverse image of $D$ under the mapping $f$ as $f^{-1}(D)=D\setminus\Gamma$. Since $D$ necessarily includes all vertices from $\Gamma$, it excludes any vertices from $N_{\scriptscriptstyle C}(\Gamma)$. Therefore, $f^{-1}(D)$ is clearly a subset of $V(C\setminus\Gamma)$ that contains all vertices in $\Theta$, while simultaneously avoiding all vertices in $\Theta^{\mathsf{c}}\setminus\Gamma$ and $N_{\scriptscriptstyle C}(\Gamma)$.

### v4

# A Study On The Graph Formulation Of Union Closed Sets Conjecture

@@ Abstract
The Union Closed Sets Conjecture is one of the most renowned problems in combinatorics. Its appeal lies in the simplicity of its statement contrasted with the potential complexity of its resolution. The conjecture posits that, in any union closed family of sets, there exists at least one element that appears in at least half of the sets within the family.
@@
Keywords: Union Closed Sets Conjecture, Frequency, Family of sets, Member sets, Pendant vertices.
@@ 1 Introduction
The Union Closed Sets Conjecture (UCC), also known as Frankl’s Conjecture, was introduced by Peter Frankl in 1979. This conjecture pertains to families of sets where the union of any two member sets remains a member of the family, called union closed families. For a given union closed family $\mathcal{F}$, the universe $U(\mathcal{F})$ is the union of all its member sets, and the frequency of an element $x\in U(\mathcal{F})$ is the count of member sets containing $x$. The conjecture asserts that for every finite union closed family $\mathcal{F}$ ($\neq\{\emptyset\}$), there exists an abundant element, defined as an element whose frequency is at least $\frac{|\mathcal{F}|}{2}$. A family is said to satisfy the UCC condition if it contains an abundant element.
@@
    When the family includes a set of cardinality 1 or 2. [6, 7]
@@
In 2013, Bruhn et al. introduced a graph-theoretic approach to the Union Closed Sets Conjecture (UCC) [2]. A subset of vertices in a graph is termed stable if no two vertices in it are adjacent. A stable set is maximal if no additional vertices can be added without losing this property.
Conjecture 1.1 (Graph formulation of UCC).
For any bipartite graph with at least one edge, there exists a vertex in each bipartite class that lies in at most half of the maximal stable sets.
A vertex as described in Conjecture 1.1 is termed rare. We say that a bipartite graph satisfies the UCC condition if it contains a rare vertex in each of its bipartite classes. This formulation is the dual of the set version of UCC, replacing an element appearing in many sets with a vertex appearing in few maximal stable sets.
In Section 2, we establish a comprehensive connection between the set-based and graph-based formulations of the conjecture. Using this connection, we present the graph version of the known results as a natural consequence. In Section 3, we study the enumeration of maximal stable sets through graph decompositions, leading to a general theorem that provides new insights into the UCC and establishes its validity for several additional graph classes. Finally, in Section 4, we outline potential directions for future research and discuss open problems motivated by our findings.
@@ 2 Connecting The Two Formulations
The union closed family generated by a family $\mathcal{F}$, denoted $\langle\mathcal{F}\rangle$, consists of the collection of all unions of its member sets, including the empty set $\emptyset$. For a graph, the set of neighbors of a vertex subset $P$ is denoted by $N(P)$. If $P=\{x\}$ is a singleton, we use $N(x)$ for convenience [^1] When multiple graphs are involved, we use $N_{\scriptscriptstyle G}(P)$, and $N_{\scriptscriptstyle G}(x)$ respectively, to specify the graph.. A pendant vertex is a vertex of degree 1. For a graph $G$, we denote the set of all maximal stable sets by $\mathcal{B}_{\scriptscriptstyle G}$. The cardinality of $\mathcal{B}_{\scriptscriptstyle G}$ is denoted by $w_{\scriptscriptstyle G}$, which represents the number of maximal stable sets of $G$. For any vertex $x\in V(G)$, the number of maximal stable sets containing $x$ is represented by $w_{\scriptscriptstyle G}(x)$.
Throughout this section, we will refer to the two vertex classes of the bipartite graph $G$ as $X$ and $Y$. We define the incidence graph of $\mathcal{F}$ as the bipartite graph with bipartition $X=\mathcal{F}$ and $Y=U(\mathcal{F})$, and edge set $\{(S,x):S\in\mathcal{F},\,x\in U(\mathcal{F}),\,x\in S\}$. For a bipartite graph $G(X,Y)$, we refer to the incidence family of $X$ as $\mathcal{F}^{\scriptscriptstyle X}=\{N(y):y\in Y\}$. If there are no isolated vertices [^2] We can omit isolated vertices without affecting the conjecture, since an isolated vertex lies in every maximal stable set and is therefore abundant. Its presence is redundant: removing it does not change the number of maximal stable sets or the frequency of any other vertices. in $X$, then its universe is given by $U(\mathcal{F}^{\scriptscriptstyle X})=X$. As an example, Figure 2 illustrates the incidence graph corresponding to the family $\mathcal{F}^{\scriptscriptstyle X}=\big\{\{1,2,3\},\{2\},\{2,3\},\{3,4\},\{4,5\}\big\}.$ In this graph, the vertices $2,3,$ and $4$ are rare, and correspondingly, the elements $2,3,$ and $4$ are abundant in the union closed family generated by $\mathcal{F}^{\scriptscriptstyle X}$.
The following proposition gives a direct connection between the set and graph formulations of the UCC, showing that an abundant element in $\langle\mathcal{F}^{\scriptscriptstyle X}\rangle$ corresponds exactly to a rare vertex in $X$. In [2] this is proved using the incidence graph of a union closed family. Here we instead use the incidence graph of a generating family, giving a more general result. The proof is inspired by [2], with the correspondence stated explicitly.
@@
Consider any subset of vertices from $Y$, denoted as $Y^{\prime}$. It can be observed that there exists a unique maximal stable set in $G$, denoted as $S$, such that $S\cap X=X\setminus N(Y^{\prime})$. Furthermore, every maximal stable set $S$ satisfies $S\cap X=X\setminus N(Y^{\prime})$ for some $Y^{\prime}\subseteq Y$. Hence $\{X\setminus N(Y^{\prime})\mid Y^{\prime}\subseteq Y\}=\{S\cap X\mid S\in\mathcal{B}_{\scriptscriptstyle G}\}$. It is important to note that no two maximal stable sets $S_{\scriptscriptstyle 1}$ and $S_{\scriptscriptstyle 2}$ can have $S_{\scriptscriptstyle 1}\cap X=S_{\scriptscriptstyle 2}\cap X$. Therefore, the set of maximal stable sets is in one-to-one correspondence with $\{S\cap X\mid S\in\mathcal{B}_{\scriptscriptstyle G}\}=\{X\setminus N(Y^{\prime})\mid Y^{\prime}\subseteq Y\}$. Note that, a vertex $x\in X$ is rare in $G$ if and only if it is rare in the family $\{X\setminus N(Y^{\prime})\mid Y^{\prime}\subseteq Y\}$, which holds if and only if $x$ is abundant in the family $\{N(Y^{\prime})\mid Y^{\prime}\subseteq Y\}$. Now, let $F=\{N(y^{\prime})\hskip 5.69046pt|\hskip 5.69046pty\in Y\}$, and observe that the universe $U(F)$ is nothing but $N(Y^{\prime})$. It is also evident that $\{N(Y^{\prime})\hskip 5.69046pt|\hskip 5.69046ptY^{\prime}\subseteq Y\}=\{U(F)\hskip 5.69046pt|\hskip 5.69046ptF\subseteq\mathcal{F}^{\scriptscriptstyle X}\}=\langle\mathcal{F}^{\scriptscriptstyle X}\rangle$. ∎
Figure 1: An illustration of an incidence family together with its corresponding incidence graph.
Figure 2: Illustration of a 2-layered vertex in a graph, where the vertex labeled 1 is 2-layered.
@@
  • A bipartite graph with both classes of cardinality at most 12 satisfies UCC condition. [follows immediately from the previous bullet point; cf. Result 1]
@@
  • In a twin-free bipartite graph, if $|Y|\geq\tfrac{2}{3}2^{\scriptscriptstyle|X|}$, then the graph satisfies the UCC condition. [Existence of a rare vertex in $X$ follows directly from cf. Result 4, and the existence of a rare vertex in $Y$ follows directly from the previous bullet point.]
  • If a bipartite graph $G$ contains a pendant vertex, then its neighbor is rare in $G$. Additionally, if $G$ contains a vertex of degree 2, then there exists a rare vertex in the other bipartite class. [Observed in [2]; also follows from the set version, cf. Result 5]
  • Let $\mathcal{F}$ be a family in which every member set has a maximum size of 3, and each element within the sets has a frequency that does not exceed 3. Then $\langle\mathcal{F}\rangle$ satisfies UCC condition. [cf. subcubic graphs satisfies UCC condition[2]]
@@
  • Let $G$ be a bipartite graph with a minimum degree of $\delta$. If both bipartite classes have cardinalities at most $2\delta$, then $G$ satisfies UCC condition. [Follows directly from the previous bullet point; cf. Result 6]
  • An $r$-regular bipartite graph with both of its bipartite classes having cardinality at most $2r$ satisfies UCC condition. [Consequence of the previous 2 bullet points; cf. Result 6]
In the subsequent section, our focus will be on graphs that possess pendant vertices positioned in specific locations.
@@ 3 Some New Graph Results
In this section, we focus on enumerating the number of maximal stable sets of a graph by decomposing it into smaller subgraphs. We begin with the simplest case, when the graph is disconnected, which is addressed in the following proposition. We then move to a more general setting, where the graph is decomposed into two subgraphs without common edges but whose union forms the parent graph. By analyzing the properties of their common vertices, we introduce a new approach to compute the maximal stable sets of the parent graph. Later, we present a result derived directly from this theorem, which establishes the UCC for several new classes of graphs.
@@
The total number of maximal stable sets of a graph G is the product of the numbers of maximal stable sets of its connected components. Moreover, in a disconnected bipartite graph G, a vertex is rare if and only if it is rare in the connected component to which it belongs.
@@
Suppose $P$ is a component of $G$ and set Q=GPQ=G-P. Let $S_{\scriptscriptstyle 1}$ and $S_{\scriptscriptstyle 2}$ be two maximal stable sets of $P$ and $Q$ respectively. One can notice that $S_{\scriptscriptstyle 1}\cup S_{\scriptscriptstyle 2}$ is a maximal stable set of $G$. For any maximal stable set $S$ of $G$, observe that $S\cap V(P)$ and $S\cap V(Q)$ are maximal stable sets of $P$ and $Q$ respectively. This gives the number of maximal stable sets of $G$, $w_{\scriptscriptstyle G}=w_{\scriptscriptstyle P}w_{\scriptscriptstyle Q}$. One can see that a maximal stable set $S$ of $G$ contains a vertex $x\in V(P)$ if and only if $S$ is of the form $S_{1}\cup S_{2}$ where $S_{1}$ and $S_{2}$ are maximal stable sets of $P$ and $Q$ respectively and $x\in S_{1}$. Hence, for every $x\in V(P)$, $w_{\scriptscriptstyle G}(x)=w_{\scriptscriptstyle P}(x)w_{\scriptscriptstyle Q}$.
Now, let $G$ be a bipartite graph and $r$ be a vertex in $P$. Then by definition, $r$ is rare in $P$ if and only if $w_{\scriptscriptstyle P}-2w_{\scriptscriptstyle P}(r)\geq 0$. As a result, $w_{\scriptscriptstyle G}-2w_{\scriptscriptstyle G}(r)=w_{\scriptscriptstyle P}w_{\scriptscriptstyle Q}-2w_{\scriptscriptstyle P}(r)w_{\scriptscriptstyle Q}=(w_{\scriptscriptstyle P}-2w_{\scriptscriptstyle P}(r))w_{\scriptscriptstyle Q}\geq 0,$ which shows that $r$ is rare in $G$ if and only if $r$ is rare in $P$. ∎
We will now present a more intriguing result. A decomposition of a graph $G$ is a set of edge-disjoint subgraphs $H_{1},H_{2},\ldots,H_{n}$ such that $\bigcup_{\scriptscriptstyle i=1}^{\scriptscriptstyle n}H_{i}=G$. A vertex $v\in V(G)$ is defined as 2-layered if every vertex in $N_{\scriptscriptstyle G}(v)$ has a pendant vertex adjacent to it. For instance, in Figure 2, the vertex labeled $1$ is 2-layered. The vertex $1$ has four neighbors $2,3,4,$ and $5$. Each of these four vertices has at least one pendant neighbor.
@@
Let $G$ be a graph and let $\{H,I\}$ be a decomposition of $G$. Suppose that the vertices in $V(H)\cap V(I)$ form a stable set and are all $2$-layered. Then the total number of maximal stable sets of $G$ is the product of the numbers of maximal stable sets of $H$ and $I$.
As a consequence of Theorem 3.2, we obtain the following theorem, which establishes some classes of graphs for which UCC condition holds.
Theorem 3.3.
Let $G$ be a bipartite graph and $\{H,I\}$ a decomposition of $G$. Suppose the vertices in $V(H)\cap V(I)$ belong to the same bipartite class of $G$ and are all 2-layered. Then, all rare vertices in $H$ or $I$ will remain rare in $G$.
Before presenting the proofs of the theorems, we first establish some notation and lemmas. The following notation will be used throughout the section. We define $n]=\{i\in\mathbb{N}\mid i\leq n\}$, and for any subset $\Theta\subseteq[n]$, we denote its complement by $\Theta^{\mathsf{c}}=[n]\setminus\Theta$. Recall the definitions of the set of all maximal stable sets $\mathcal{B}_{G}$ and its cardinality $w_{G}$. Here, we will slightly generalize these definitions. For any vertex subsets $P,Q\subseteq V(G)$ and any subgraph $K$ of $G$, the notation $\mathcal{B}_{\scriptscriptstyle K}(P,\overline{Q})$ denotes the set of all maximal stable sets of $K$ that include all vertices in $P$ and exclude all vertices in $Q$. The cardinality of $\mathcal{B}_{\scriptscriptstyle K}(P,\overline{Q})$ is denoted by $w_{\scriptscriptstyle K}(P,\overline{Q})$. We write $K\setminus P$ to denote the subgraph obtained by removing the vertices in $P$ and all edges incident to them from $K$.
Figure 3: Illustration of Lemma 3.4 with graph $H$, stable set $[n]$ consisting of all 2-layered vertices, and the subset $\Gamma\subseteq\Theta^{\mathsf{c}}$.
Figure 4: An illustration of Lemma 3.6, showing the decomposition of graph $G$ into subgraphs $H$ and $I$. The common vertex set $[n]$ is highlighted together with its partition into $\Theta$, $\Theta^{\mathsf{c}}$, $\Gamma\subseteq\Theta^{\mathsf{c}}$, and $\Psi\subseteq\Theta^{\mathsf{c}}\setminus\Gamma$.
@@
Let $H$ be a graph with $n]\subseteq V(H)$ such that $[n]$ is a stable set and all vertices of $[n]$ are 2-layered. For any $\Theta\subseteq[n]$, $\Gamma\subseteq\Theta^{\mathsf{c}}$, and $a\in V(H)\setminus\Gamma$, the following identities hold:
$w_{\scriptscriptstyle H\setminus\Gamma}(\Theta,\overline{\Theta^{\mathsf{c}}\cup N_{\scriptscriptstyle H}(\Gamma)})=w_{\scriptscriptstyle H}(\Theta\cup\Gamma,\overline{\Theta^{\mathsf{c}}\setminus\Gamma})$ (1)
$w_{\scriptscriptstyle H\setminus\Gamma}(\{a\}\cup\Theta,\overline{\Theta^{\mathsf{c}}\cup N_{\scriptscriptstyle H}(\Gamma)})=w_{\scriptscriptstyle H}(\{a\}\cup\Theta\cup\Gamma,\overline{\Theta^{\mathsf{c}}\setminus\Gamma})$ (2)
@@
Consider the mapping $f:\mathcal{B}_{\scriptscriptstyle H\setminus\Gamma}(\Theta,\overline{\Theta^{\mathsf{c}}\cup N_{\scriptscriptstyle H}(\Gamma)})\to\mathcal{B}_{\scriptscriptstyle H}(\Theta\cup\Gamma,\overline{\Theta^{\mathsf{c}}\setminus\Gamma})$ defined by $f(B)=B\cup\Gamma$. Let $B\in\mathcal{B}{\scriptscriptstyle H\setminus\Gamma}(\Theta,\overline{\Theta^{\mathsf{c}}\cup N{\scriptscriptstyle H}(\Gamma)})$, observe that $f(B)$ includes all vertices from $\Theta\cup\Gamma$ and excludes all vertices from $\Theta^{\mathsf{c}}\setminus\Gamma$. Since $B$ is a stable set in $H\setminus\Gamma$ and contains no vertices from $N_{\scriptscriptstyle H}(\Gamma)$, it follows that $f(B)$ is stable in $H$. Assume, for contradiction, that $f(B)$ is not a maximal stable set in $H$. Then, there exists a vertex $v\in V(H)$ such that $v\notin f(B)$ and $B\cup\Gamma\cup\{v\}$ is stable in $H$. This implies that $B\cup\{v\}$ is stable in $H\setminus\Gamma$, contradicting the maximality of $B$ in $H\setminus\Gamma$. Hence, $f(B)\in\mathcal{B}_{\scriptscriptstyle H}(\Theta\cup\Gamma,\overline{\Theta^{\mathsf{c}}\setminus\Gamma})$.
@@
Next, consider a set $D\in\mathcal{B}_{\scriptscriptstyle H}(\Theta\cup\Gamma,\overline{\Theta^{\mathsf{c}}\setminus\Gamma})$. We define the inverse image of $D$ under the mapping $f$ as $f^{-1}(D)=D\setminus\Gamma$. Since $D$ necessarily includes all vertices from $\Gamma$, it excludes any vertices from $N_{\scriptscriptstyle H}(\Gamma)$. Therefore, $f^{-1}(D)$ is clearly a subset of $V(H\setminus\Gamma)$ that contains all vertices in $\Theta$, while simultaneously avoiding all vertices in $\Theta^{\mathsf{c}}\setminus\Gamma$ and $N_{\scriptscriptstyle H}(\Gamma)$.